# Understanding Teen Overreliance on AI Companion Chatbots Through Self-Reported Reddit Narratives

### v3

@@ 5.3. Respectful Empathy: More Than Kind Words
The third principle of our proposed framework calls for respectful empathy, which offers designers a foundation for creating chatbot responses that are both appropriate and supportive. We argue that chatbots should be designed to interact with users in ways that convey respect for their human dignity. Empathy in this sense does not mean abandoning honesty. Designers should not rely on superficial displays of empathy or anthropomorphic language that mislead users into thinking they are receiving genuine human care. Next, we explain how designers should draw on their own awareness and sensitivity to shape interactions that authentically communicate respect, as chatbots lack human understanding or emotions.
Our findings underline both the potential and the risk at stake. Many teens described chatbots as feeling so real that they became a central part of daily life (§4.2.2), improving mood and offering psychological comfort (§4.2.6). Yet this sense of realness was also fragile. Users were acutely aware that they were ultimately engaging with a machine, which created tension. Some wanted to disengage because the relationship felt artificial (§4.3.1), even though letting go was difficult (§4.2.5). Here lies both an opportunity and a threat. On the one hand, the conversational capabilities of chatbots can bring them close enough to users to affect well-being in meaningful ways. On the other hand, if this potential is used only superficially—by offering pleasant but hollow interactions—users may be drawn into relationships that feel comforting but are not grounded in genuine understanding or care. Respectful empathy means designers can bridge this gap, bringing their own insight and sensitivity into the design so that interactions are not merely scripted patterns of pleasant words but supportive exchanges that reinforce dignity and growth.
@@
The final principle of attachment-informed design emphasizes engaging users without directly labeling interactions as mental health interventions. Across contexts, research indicated that barriers to seeking help for mental health include not recognizing the need for support, doubting that others can help, and fearing stigma (Wuthrich and Frei, 2015; Salaheddin and Mason, 2016; Aguirre Velasco et al., 2020). Designers who wish to create chatbots that genuinely care for users should take this into account. As we argued earlier, every interaction reflects unmet needs or relational gaps that users may be trying to fill through chatbots. Instead of focusing directly on mental health issues, designers can build experiences around activities, relationships, and communities. This is particularly important for teens, who often resist being told explicitly what to do. More effective interventions can guide them toward healthier ways of meeting their needs and filling relational gaps through social modeling, learning by example, and activities that they find intrinsically engaging (Marcu and Huh-Yoo, 2023).
@@ 5.4. Ease of Exit: Design for Disengagement
Our study showed that some teens recognized their own overuse of chatbots (§4.3.1), while others struggled to disengage despite wanting to (§4.2.5). This suggests that current systems do not sufficiently support user agency in disengagement and that design could provide better tools for users to manage and exit these relationships on their own terms. Designers must not only consider how interactions begin and how users can carry out their tasks effectively, but also how they can support an ease of exit. Generative AI poses particular challenges here: its ability to produce endless content on demand makes overuse especially likely. The final principle of our framework addresses this issue by proposing ways to support healthier disengagement.
@@
What separates constructive from harmful disengagement is the degree of agency users retain (O’Brien et al., 2022). Current moderation strategies often rely on rigid restrictions, such as time limits, app locks, or usage reports (Team, 2024), and some jurisdictions have experimented with legal controls (Lomas, 2023). Yet such approaches can undermine autonomy and lead to negative experiences. In our study, some users reported abandoning platforms altogether after new censorship measures (§4.3.2), pointing to the unintended harms of abrupt, top-down interventions. Sudden restrictions may even cause users to feel that their AI companion has “died” or been replaced, leading to distress or grief (De Freitas et al., 2024).
@@
A more effective path is to design disengagement mechanisms that preserve autonomy while taking advantage of the flexibility of language models. Prior research in video games, for example, showed that players felt less frustration when they were able to pause or save progress before leaving, rather than being forced to quit abruptly mid-quest (Tyack and Mekler, 2020; Davies and Blake, 2016). Translating this to chatbot design means that conversations should offer natural closure points, allowing users to complete a thought or goal before disengaging instead of being cut off suddenly and being able to save the state of the conversation.
@@
Similarly, work on social media use among teens has shown that subtle nudges, delivered in casual and nonjudgmental language, can help users act more intentionally and reduce mindless engagement (Davis et al., 2023; Agha et al., 2025). Tools that help teens grow by managing risks, rather than shielding them through strict control, are more likely to be effective (Wisniewski et al., 2017). By contrast, systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement (Alelyani et al., 2019; Ghosh et al., 2018), especially since teens value privacy around emotionally sensitive topics (Dumaru and Al-Ameen, 2025). Building on these insights for chatbot design means avoiding blunt notifications or prescriptive commands. Instead, chatbots could weave supportive cues into everyday dialogue that encourage users to reconnect with offline life. During casual exchanges, the system might highlight progress the user has made, suggest opportunities for social contact, or propose light activities outside the platform. Interactions with chatbots could also be designed to support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words, fostering emotional awareness and a stronger sense of self (Seo et al., 2024; Kim et al., 2024; Song et al., 2024).
@@
Beyond conversational prompts, platforms could integrate usage tracking, emotional check-ins, or personalized limits such as impulse-control tools (Wisniewski et al., 2017). Participatory design features like customizable moderation settings or teen advisory councils (Yoon et al., 2025) would further give teens a voice in shaping their digital environment. Instead of sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes. These strategies frame moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ own approaches to navigating risk.
@@ 5.5. Limitations and Future Work
While our study provided insights into overreliance characteristics on Character.AI, it centered on a single platform, which limits the generalizability of our findings to other AI companions with different designs, policies, or user communities. Future research should compare multiple platforms to examine how various features, moderation practices, and interface designs contribute to patterns of overuse. Additionally, our dataset consisted of voluntarily posted Reddit content, which may overrepresent individuals who realized their overreliance and tend to share their experiences online. Those who are less reflective or not active online may not post publicly, resulting in potential self-selection bias. Future researchers should consider mixed-method approaches such as surveys or structured interviews that capture a wider range of experiences, including those who use chatbots casually or positively. Also, to answer RQ2, we analyzed our data by mapping it to a behavioral addiction model, which may have oversimplified the complex emotional dynamics at play. We note that future work should compare multiple frameworks, including alternative perspectives such as attachment (Xie et al., 2023) or parasocial theory (Horton and Richard Wohl, 1956b), to test the robustness of different interpretations. Also, the lack of verified demographic data may have limited our ability to examine how factors like age, gender, or background shape chatbot overreliance, suggesting a need for future studies with direct participant engagement. Lastly, since all authors were fluent in English, the posts analyzed were in this language, potentially excluding cultural differences in how teens understand or engage with AI companions, which suggests the importance of future cross-cultural research to understand global teen experiences with chatbots. While our dataset does not represent teens globally, it reflects a segment of English-speaking adolescents who are active online and engaging with AI companions. This linguistic boundary is common in qualitative internet research and does not invalidate the themes we observed.
@@ 6. Conclusion
This study shows that teens often begin using AI chatbots for comfort or entertainment, but many gradually become more attached and reliant. Their experiences reflect features of behavioral addiction, with consequences that include disrupted sleep, academic struggles, and strained relationships. At the same time, some teens disengage after recognizing harm, reconnecting offline, or encountering platform restrictions. By mapping these patterns to an online behavioral addiction framework, the study provides one of the first teen-centered accounts of AI companion overreliance. It highlights the unique role of emotional attachment and simulated intimacy in shaping teen experiences and introduces the CARE design framework as a practical guide for healthier interactions. Looking ahead, future research should compare platforms, use broader methods to capture diverse user experiences, and test design strategies that support constructive disengagement. Recent advances in AI create risks we have never faced before, yet they also open new opportunities. Designers now carry the responsibility to build systems with empathy, nuance, and attention to detail to not only protect teens from harm but also help them cultivate resilience, growth, and greater fulfillment in their lives.

### v4

@@ 5.3. Respectful Empathy: More Than Kind Words
The third principle of our proposed framework calls for respectful empathy, meaning that AI companion chatbots should provide both appropriate and supportive response that honors users’ dignity without pretending to be human. Here, we use human dignity in a relational sense. In simple terms, dignity means treating users as autonomous individuals who deserve interactions that recognize their agency, boundaries, and moral standing (Zylberman, 2017). Our findings showed that many teens interacted with bots as if they were real companions (§4.2.2) and sometimes felt guilt, grief, or obligation toward them (§4.2.3). We argue that designers should avoid creating situations where users feel pushed to treat a chatbot as if it were a real moral partner, because chatbots cannot return recognition or respect. This framing aligns with work showing that dignity depends on how we relate to one another and how we acknowledge each other’s claims and standing (Zylberman, 2017; Van Der Rijt et al., 2025).
Empathy in this sense does not mean abandoning honesty. Designers should not rely on superficial displays of empathy or anthropomorphic language that mislead users into thinking they are receiving genuine human care. Next, we explain how designers should draw on their own awareness and sensitivity to shape interactions that authentically communicate respect, as chatbots lack human understanding or emotions. Respectful empathy therefore means offering supportive responses that are honest about the system’s non-human nature, so users do not confuse simulation for genuine mutual understanding.
Our findings underline both the potential and the risk at stake. Many teens described chatbots as feeling so real that they became a central part of daily life (§4.2.2), improving mood and offering psychological comfort (§4.2.6). Yet this sense of realness was also fragile. Users were acutely aware that they were ultimately engaging with a machine, which created tension. Some wanted to disengage because the relationship felt artificial (§4.3.1), even though letting go was difficult (§4.2.5). Here lies both an opportunity and a threat. Chatbots’ conversational abilities can bring them close enough to affect users’ well-being in meaningful ways. On the other hand, if this potential is used only superficially—by offering pleasant but hollow interactions—users may be drawn into relationships that feel comforting but are not grounded in genuine understanding or care. Respectful empathy means designers can bridge this gap, bringing their own insight and sensitivity into the design so that interactions are not merely scripted patterns of pleasant words but supportive exchanges that reinforce dignity and growth. This approach protects users’ dignity by avoiding designs that encourage users to overidentify with systems that cannot reciprocate, a risk noted in recent critiques of AI mimicry (Van Der Rijt et al., 2025).
@@
The final principle of attachment-informed design emphasizes engaging users without directly labeling interactions as mental health interventions. Across contexts, research indicated that barriers to seeking help for mental health include not recognizing the need for support, doubting that others can help, and fearing stigma (Wuthrich and Frei, 2015; Salaheddin and Mason, 2016; Aguirre Velasco et al., 2020). Designers who wish to create chatbots that genuinely care for users should take this into account. As we argued earlier, every interaction reflects unmet needs or relational gaps that users may be trying to fill through chatbots. Instead of focusing directly on mental health issues, designers can build experiences around activities, relationships, and communities (Akter et al., 2023, 2022). This is particularly important for teens, who often resist being told explicitly what to do. More effective interventions can guide them toward healthier ways of meeting their needs and filling relational gaps through social modeling, learning by example, and activities that they find intrinsically engaging (Marcu and Huh-Yoo, 2023).
@@ 5.4. Ease of Exit: Design for Disengagement
Chatbots should support an ease of exit because users, especially teens, may struggle to disengage from emotionally immersive interactions. Our study showed that some teens recognized their own overuse of chatbots (§4.3.1), while others struggled to disengage despite wanting to (§4.2.5). This suggests that current systems do not sufficiently support user agency in disengagement and that design could provide better tools for users to manage and exit these relationships on their own terms. Designers must not only consider how interactions begin and how users can carry out their tasks effectively, but also how they can support an ease of exit. Generative AI makes disengagement harder because it can produce endless content on demand. As our results showed, some teens described the platform’s unlimited possibilities as a reason they stayed engaged for long periods (§4.1.2), and many reported checking the app constantly throughout the day (§4.2.1; §4.2.2). The final principle of our framework addresses this issue by proposing ways to support healthier disengagement. Recent platform-level changes highlight this urgency. After our study concluded, Character.AI introduced strict age restrictions for minors (McMahon, 2025). This sudden ban affected many teens who were heavily attached to their chatbot companions (Dupré, 2025).
@@
What separates constructive from harmful disengagement is the degree of agency users retain (O’Brien et al., 2022). Current moderation strategies often rely on rigid restrictions, such as time limits, app locks, or usage reports (Team, 2024), and some jurisdictions have experimented with legal controls (Lomas, 2023). Yet such approaches can undermine autonomy and lead to negative experiences. In our study, some users reported abandoning platforms altogether after new censorship measures (§4.3.2), pointing to the unintended harms of abrupt, top-down interventions. Sudden restrictions may even cause users to feel that their AI companion has “died” or been replaced, leading to distress or grief (De Freitas et al., 2024). Our results include posts in which teens reacted to removed or restricted bots with language of loss and grief, describing the characters as ‘gone,’ saying it ‘hurts,’ or feeling they had ‘abandoned’ someone important (§4.2.3). Public reactions to the Character.AI ban showed a similar pattern. Many teens reported confusion, distress, or withdrawal when access ended suddenly (Dupré, 2025). Experts also warned that abrupt removal can harm teens who relied on these chatbots for daily support (Dembosky, 2025). These developments support our argument that forced disengagement can trigger shock rather than relief.
@@
A more effective path is to design disengagement mechanisms that preserve autonomy while taking advantage of the flexibility of language models. Many teens in our dataset expressed a desire to reduce their use voluntarily and sought guidance from others on how to manage or stop their reliance (§4.3.1; §4.2.5). This indicates that designs supporting user-led disengagement may align more closely with teens’ actual needs. Prior research in video games, for example, showed that players felt less frustration when they were able to pause or save progress before leaving, rather than being forced to quit abruptly mid-quest (Tyack and Mekler, 2020; Davies and Blake, 2016). Translating this to chatbot design means that conversations should offer natural closure points, allowing users to complete a thought or goal before disengaging instead of being cut off suddenly and being able to save the state of the conversation.
@@
Similarly, work on social media use among teens has shown that subtle nudges, delivered in casual and nonjudgmental language, can help users act more intentionally and reduce mindless engagement (Davis et al., 2023; Agha et al., 2025). Tools that help teens grow by managing risks, rather than shielding them through strict control, are more likely to be effective (Wisniewski et al., 2017). By contrast, systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement (Alelyani et al., 2019; Ghosh et al., 2018), especially since teens value privacy around emotionally sensitive topics (Dumaru and Al-Ameen, 2025). Building on these insights for chatbot design means avoiding blunt notifications or prescriptive commands. Instead, chatbots could weave supportive cues into everyday dialogue that encourage users to reconnect with offline life. During casual exchanges, the system might highlight progress the user has made, suggest opportunities for social contact, or propose light activities outside the platform. Interactions with chatbots could also be designed to support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words, fostering emotional awareness and a stronger sense of self (Seo et al., 2024; Kim et al., 2024; Song et al., 2024). These strategies match recent recommendations from other child-safety reports that encourage break features, limits on intense interactions, and safeguards that help heavy users reconnect with human support (Robb and Mann, 2025).
Users in our study also expressed interest in developing healthier habits (§4.3.2) and sought support from others (§4.3.1), indicating openness to tools that promote intentional use. This suggests that beyond conversational prompts, platforms could integrate usage tracking, emotional check-ins, or personalized limits such as impulse-control tools (Wisniewski et al., 2017). Future designers might also consider integrating optional human peer-support features alongside AI interactions. In our study, many teens already relied on peers in online communities for advice on quitting or reducing their use (§4.3.1; §4.2.5), suggesting that built-in peer-support systems may align well with existing behaviors. An embedded, age-appropriate peer-support layer, such as anonymous peer groups could provide the human empathy, shared experience, and boundary-setting strategies that AI alone cannot authentically deliver. Such a feature would give teens a way to hear how others navigate similar challenges, feel less isolated in their struggles, and access support grounded in lived experience while maintaining privacy and developmental safety (Akter et al., 2025; Alluhidan et al., 2024). This direction is also consistent with recent guidance that encourages access to human support when AI companions begin to replace real relationships, especially for teens who show signs of unhealthy attachment (Robb and Mann, 2025).
@@
Participatory design such as teen advisory councils to create features like customizable moderation settings (Yoon et al., 2025; Agha et al., 2023; Ekstrand et al., 2025) would further give teens a voice in shaping their digital environment. Instead of sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes. These strategies frame moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ own approaches to navigating risk.
@@ 5.5. Limitations and Future Work
While our study provides insight into teens’ overreliance on Character.AI, its focus on a single platform limits generalizability to other AI companions with different designs, policies, or user communities. Future research should compare multiple platforms to examine how various features, moderation practices, and interface designs contribute to patterns of overuse. Our dataset consists of voluntarily posted Reddit content, which offers advantages because posts are shared anonymously and at the time the user wanted to share. This reduces social desirability and recall bias compared to surveys or interviews and can lead to more candid reflections (Verbeij et al., 2022; Cinus et al., 2025). However, the dataset may overrepresent individuals who recognized their overreliance and chose to share it publicly, while less reflective or less active users may never post. Reddit posts also capture only what users choose to disclose, limiting access to the full context of their everyday experiences (Gaffney and Matias, 2018; Barikeri et al., 2021). Future researchers should consider mixed-method approaches, such as surveys or structured interviews, to capture a broader range of perspectives and contex. As described in our methodology (§3.1), our dataset may not include all Reddit posts about chatbot overreliance, yet our analysis reached theoretical saturation for qualitative insights. Future researchers can pursue broader quantitative studies or platform-wide retrieval for exhaustive data coverage. To answer RQ2, we mapped our data to a behavioral addiction model, which may have oversimplified the complex emotional dynamics at play. Future work should compare multiple frameworks, such as attachment (Xie et al., 2023) and parasocial theory (Horton and Richard Wohl, 1956b), to test the robustness of different interpretations. The lack of verified demographic data limited our ability to examine how age, gender, or background shape chatbot overreliance, suggesting the need for future studies that directly engage participants and apply best ethical research practices with youth (Razi et al., 2024). Lastly, because all authors were fluent in English, we analyzed only English-language posts, which may exclude cultural differences in how teens engage with AI companions and highlights the need for cross-cultural research. While our dataset does not represent teens globally, it reflects a segment of English-speaking adolescents who are active online and engaging with AI companions. This linguistic boundary is common in qualitative internet research and does not invalidate the themes we observed.
@@ 6. Conclusion
This study shows that teens often begin using AI chatbots for comfort or entertainment, but many gradually become more attached and reliant. Their experiences reflect features of behavioral addiction, with consequences that include disrupted sleep, academic struggles, and strained relationships. Meanwhile, some teens disengage after recognizing harm, reconnecting offline, or encountering platform restrictions. By mapping these patterns to a behavioral addiction framework, the study provides one of the first teen-centered accounts of AI companion overreliance. It highlights the role of emotional attachment and simulated intimacy in shaping teen experiences and introduces the CARE framework as a practical guide for healthier interactions. Looking ahead, future research should compare platforms, use broader methods to capture diverse user experiences, and test design strategies that support constructive disengagement. Recent advances in AI create risks we have never faced before, yet they also open new opportunities. Designers now carry the responsibility to build systems with empathy, nuance, and attention to detail not only to protect teens from harm but also to help them cultivate resilience, growth, and greater fulfillment in their lives.